# Authoritarian Recursions: How Fiction, History, and AI Reinforce Control in Education, Warfare, and Discourse

### v2

@@ Abstract
This article introduces the concept of authoritarian recursion to describe how artificial intelligence (AI) systems increasingly mediate control across education, warfare, and digital discourse. Drawing on critical discourse analysis and sociotechnical theory, the study reveals how AI-driven platforms delegate judgment to algorithmic processes, normalize opacity, and recursively reinforce behavioral norms under the guise of neutrality and optimization. Case studies include generative AI models in classroom surveillance, autonomous targeting in military AI systems, and content curation logics in platform governance.
Rather than treating these domains as disparate, the paper maps their structural convergence within recursive architectures of abstraction, surveillance, and classification. These feedback systems do not simply automate tasks—they encode modes of epistemic authority that disperse accountability while intensifying political asymmetries. Through cultural and policy analysis, the article argues that authoritarian recursion operates as a hybrid logic, fusing technical abstraction with state and market imperatives. The paper concludes by outlining implications for democratic legitimacy, human oversight, and the political design of AI governance frameworks.
This framework contributes to emerging debates on algorithmic accountability by foregrounding how recursion acts not merely as a technical function but as a sociopolitical instrument of control.
Keywords: authoritarian recursion, AI governance, algorithmic accountability, surveillance, recursive control, critical discourse analysis, sociotechnical systems, educational technology, military AI, platform regulation
Introduction
Artificial intelligence (AI) is often described in terms of promise—optimization, personalization, predictive precision—but it is equally a technology of recursion. It does not merely operate within existing systems; it loops back upon them, amplifying their logics, stabilizing their asymmetries, and obscuring their origins. This paper names this recursive mechanism authoritarian recursion: the systematic reinforcement and evolution of control structures through algorithmic automation.
Authoritarian recursion is not a metaphor. It describes how AI systems—deployed in warfare, education, and content governance—codify prior patterns of judgment, visibility, and exclusion. These systems are rarely neutral. Their architectures inherit militarized epistemologies, bureaucratic incentives, and market logics that prioritize efficiency over deliberation, surveillance over solidarity. What emerges is not a clean break with the past, but a patterned automation of its most coercive functions.
This argument unfolds across three interrelated sites: drone warfare, algorithmic proctoring, and content moderation. These domains are not chosen for breadth but for depth—they reveal how AI mediates moral judgment under conditions of scale, uncertainty, and institutional inertia. Far from discrete sectors, they form a recursive matrix in which power circulates and legitimates itself through code.
The analysis proceeds interpretively, drawing on critical media theory, relational ethics, and fictional narratives that illuminate the epistemic stakes of automation. Cultural texts like Black Mirror, The Terminator, and Nineteen Eighty-Four function not as speculative detours but as epistemological tools. They dramatize how automation, once delegated, escapes retraction—how decisions made by machines are increasingly insulated from human redress. As Cave and Dihal argue, such representations do not merely reflect public fears; they actively shape institutional responses to AI [1].
This inquiry also grounds itself in contemporary institutional discourse. Corporate white papers, governmental defense protocols, and educational AI policies offer insight into how algorithmic systems are rationalized and normalized. Reports such as Meta’s transparency roadmap [2], Google’s responsible AI framework [3], and the Global Network Initiative’s human rights standards [4] reveal not only what these actors claim AI can do, but also what they believe it ought to do—and for whom.
AI is not deployed in a political vacuum. It is embedded within governance architectures that frame ethical responsibilities, determine thresholds of acceptability, and allocate moral authority to technical systems. To understand this embedding is not to reject technology but to interrogate its premises: Who defines fairness? Who delegates oversight? Who becomes invisible under the regime of the algorithm?
This paper argues that authoritarian recursion is not an aberration but a structural tendency—a design pattern disguised as inevitability. Recognizing it demands we move beyond interface-level reforms and ask harder questions about institutional legitimacy, recursive design, and the politics of automation itself.
@@ 1 Literature Review
This section reviews the scholarly and historical terrain across three domains where AI technologies operationalize control: military automation, educational surveillance, and algorithmic curation in digital discourse. While each context has sector-specific concerns, they share structural features—namely, opacity, delegation of judgment, and recursive feedback—that embed normative authority into automated systems.
@@ 1.1 Military AI and the Automation of Violence
These concerns align with the Fairness, Accountability, and Transparency (FAccT) framework, which argues that fairness in automated decision-making must extend beyond output metrics to include contextual sensitivity, procedural redress, and meaningful oversight [6, 7]. In military AI, accountability often becomes nominal—tethered to protocols rather than substantive ethical reflection.
Historical precedents further complicate the notion of technological neutrality. During World War II, IBM’s punch card infrastructure was deployed by Nazi Germany for logistics and census operations—systems later used to facilitate genocide [8]. These tools encoded forms of bureaucratic rationality eerily analogous to today’s algorithmic architectures. As Asaro argues, the distancing of lethal decisions from moral responsibility through automation represents a dangerous ethical shift [9].
Speculative fiction reinforces this critique. Narratives like The Terminator, RoboCop, and Black Mirror episodes offer dystopian imaginaries of militarized AI. According to Cave et al., “the future imagined in fiction is often realized not because it is inevitable, but because it is ideologically compatible with dominant institutions” [10, p. 75]. These cultural texts anticipate how automation narratives rationalize political authority and normalize autonomous violence.
@@ 1.2 Educational AI: Automation of Surveillance and Discipline
Here, too, the FAccT triad is often invoked as a remedy. However, operational deployments rarely meet its normative thresholds. Mittelstadt et al. emphasize that fairness requires attention to context and historical inequalities, not just statistical parity [6]. Yet student-users typically lack access to the internal logic of these systems and have little recourse to challenge their outputs.
Noble argues that “algorithmic decision systems often act as new instruments of racial and economic profiling” [12, p. 34]. This echoes Selwyn’s concerns that digital surveillance in education creates a system “in which suspicion is automated and dissent is pathologized” [13]. Such platforms reproduce the logic of panoptic discipline, as theorized by Foucault, where constant surveillance internalizes conformity [14].
Moreover, the political values behind these technologies often go unquestioned. Gilliard and Selwyn contend that “continued adoption of proctoring technologies in public education exposes a fundamental clash of politics,” where commercial priorities of security and efficiency override pedagogical values of equity and trust [15, p. 197]. These systems operationalize discipline not through direct coercion, but through the automation of suspicion and reduction of students to behavioral data.
@@ 1.3 AI in Discourse and Propaganda: Curation as Control
Gillespie emphasizes that platforms are not neutral hosts but “custodians of public discourse” who shape access to visibility through inscrutable recommendation logics [17, p. 197]. Zuboff describes this shift as “instrumentarian power,” wherein behavior is not repressed but tuned through predictive analytics and behavioral nudging [18, p. 377].
@@
Together, these patterns constitute what may be termed an authoritarian recursion—a self-reinforcing cycle in which AI technologies encode, naturalize, and propagate control logics across domains. As behavior is shaped by algorithmic outputs, those outputs become future inputs, creating recursive systems that reinforce the very assumptions they were built upon.
This synthesis lays the foundation for the case study analysis that follows, which further illustrates how intelligent systems materialize these recursive dynamics in real-world governance structures.
@@ 2 Methodology
The analysis proceeds from an interpretive, critical-theoretical perspective, aimed at understanding how artificial intelligence (AI) systems participate in the reproduction of normative control across different social domains. Rather than offering empirical generalizations or causal inferences, the approach foregrounds how meaning is made, contested, and stabilized through discourse, design narratives, and institutional practice.
At the core of this inquiry lies Critical Discourse Analysis (CDA), particularly as theorized by Norman Fairclough. He contends that “discourse is both constitutive and constituted; it contributes to the shaping of social structures while being shaped by them” [22]. In other words, how we talk about AI—whether as intelligent, efficient, autonomous, or fair—feeds back into the very architectures we build. CDA enables a focus on these feedback mechanisms: where language meets infrastructure, and where discourse becomes design.
The analysis draws on three main types of material: theoretical literature, cultural narratives, and institutional documents. First, foundational texts in AI ethics, surveillance capitalism, and platform governance provide conceptual grounding. These include works by Zuboff, Coeckelbergh, and Gillespie, whose insights into autonomy, control, and datafication shape the framing of recursive algorithmic systems [18, 23, 17].
Second, cultural narratives—especially speculative fiction such as Orwell’s Nineteen Eighty-Four [24], James Cameron’s The Terminator [25], and episodes from Charlie Brooker’s anthology series Black Mirror [26]—are employed not as empirical data, but as heuristic provocations. These stories surface emerging ethical dilemmas in ways that formal models often overlook. As Cave and Dihal argue, cultural imaginaries help structure how societies anticipate and evaluate technologies before they fully arrive [1].
Third, the analysis incorporates public-facing and institutional documents, including educational policies on algorithmic proctoring [13], defense white papers on autonomous weapons systems [9], and corporate reports outlining the principles and practices of algorithmic content moderation. For instance, transparency initiatives like Meta’s system cards [2], Google’s Responsible AI framework [3], and the Global Network Initiative’s Principles on Freedom of Expression and Privacy [4] reveal how AI ethics is being operationalized at scale. These materials allow the study to trace not only how AI systems are designed and deployed, but how their legitimacy is framed through policy discourse and institutional self-regulation.
What holds these sources together is not empirical coherence but thematic resonance. All engage with questions of visibility, judgment, autonomy, and legitimacy—questions that lie at the center of both AI deployment and ethical inquiry. While some might ask whether these sources can be analyzed together, the real question is: how could they not be, given how power now flows across sectors and interfaces?
Fictional systems like Skynet, for example, are not invoked because they “predict” the future but because they illuminate a logic already at work: the automation of escalation, the erasure of moral deliberation, and the insulation of decision-making from relational accountability. Skynet is less a fantasy than a frame—one that helps articulate why human oversight cannot be treated as an afterthought in system design.
This interpretive approach has its limitations. It does not claim neutrality, nor does it offer quantifiable findings. But it does aim to clarify the stakes of a rapidly changing technological order by drawing lines between the past, the possible, and the political. The hope is that such lines might help us see AI systems not just as tools to be used, but as structures to be questioned.
@@ 3 Case Studies
AI does not act in isolation—it is embedded in systems that shape how knowledge, risk, and authority are organized. The following three domains—education, warfare, and discourse—reveal how intelligent systems delegate judgment, automate suspicion, and reinforce pre-existing norms through recursive feedback loops. These cases are not offered as an exhaustive taxonomy but as analytical vignettes into the institutional logics that AI both reflects and reconfigures.
3.1 Education: Algorithmic Proctoring and Epistemic Discipline
Remote proctoring platforms such as Proctorio and ExamSoft are marketed as scalable solutions for maintaining academic integrity in digital environments. Yet these systems pre-encode suspicion: gaze aversion, posture shifts, ambient silence—each is parsed as a probabilistic indicator of misconduct. In doing so, they invert the burden of proof and presume bad faith as a system default.
As Gilliard and Selwyn argue, such platforms enact “a shift in educational values, from those grounded in equity and social justice to those centered on control, surveillance, and suspicion” [15]. The proctor is no longer a relational figure but a predictive apparatus—opaque, unaccountable, and disembedded from context.
This dynamic echoes Foucault’s account of disciplinary power, where visibility enforces internalized control: “He who is subjected to a field of visibility… assumes responsibility for the constraints of power” [14, p. 202]. Algorithmic proctoring does not merely monitor students—it restructures the epistemology of assessment. Pedagogy becomes risk classification. Trust becomes anomaly detection. Judgment is replaced by metrics optimized for auditability. In this configuration, learning is subordinated to legibility.
3.2 Warfare: Autonomous Weapons and Delegated Lethality
Autonomous weapons systems (AWS), including platforms like Israel’s Harpy drone and the U.S. MQ-9 Reaper, are designed to identify and engage targets with minimal or no human oversight. Framed as innovations in precision warfare, they operationalize a deeper abstraction: the displacement of moral agency by machinic logics.
Peter Asaro warns that “removing humans from the process of selecting and engaging targets dehumanizes lethal decision-making and thereby undermines fundamental principles of human rights and international humanitarian law” [9, p. 696]. Crucially, this is not just a technical transition—it is a normative rupture. Asaro continues: “Delegating these decisions to machines undermines the moral responsibility of the individuals and institutions involved, and therefore threatens the integrity of both legal and moral systems of accountability” [9, p. 701]. What emerges is not efficiency, but moral evacuation. Killing becomes a system function.
Automation also reconfigures the phenomenology of violence. “The psychological distance of the operator from the act of killing has long been a concern in the use of drones, but automation increases this distance even further” [9, p. 695]. Warfare becomes unbounded by the human sensorium—executed by systems, rationalized by code, and shrouded in algorithmic opacity.
Speculative fiction offers an uncanny mirror. Skynet—the sentient AI in The Terminator—is often dismissed as fantasy. But its structure reveals real-world logics: recursive autonomy, strategic opacity, and threat generalization. Cave and Dihal note that such imaginaries “actively shape how people think about what AI is and what it might become” [1, p. 690]. Skynet does not forecast—it refracts. It models what happens when warfighting systems outpace the ethics meant to constrain them.
3.3 Discourse: Algorithmic Curation and Visibility Control
Social media platforms like TikTok, Facebook, and YouTube no longer merely host content—they algorithmically shape the public sphere. Their recommender systems optimize for engagement, modeling user behavior to predict what content will provoke further interaction. But when engagement becomes the principal metric, what happens to truth?
Gillespie characterizes this shift as a move from editorial judgment to “infrastructural moderation” [17, p. 197]. Visibility is no longer the outcome of deliberation, but the artifact of optimization. Relevance, not accuracy, drives circulation. Prior behavior feeds future exposure, forming recursive loops in which dissent is drowned out by virality.
This is not simply bias—it is infrastructural drift. As Zuboff contends, the behavioral data extracted from users becomes “raw material for a new means of behavioral control” [18, p. 377]. The platform learns what you respond to, and then governs through that responsiveness. In this feedback ecology, recommender algorithms become both mirror and motor: they reflect user behavior while simultaneously training it toward platform-compatible norms.
Here, knowledge is not curated—it is sorted. Political speech is flattened into click-through potential. Inquiry becomes indistinguishable from consumption. And in that sorting, what is sayable and knowable is quietly redefined.
Across education, warfare, and discourse, AI systems do more than optimize tasks—they reorganize ethical coordinates. They embed suspicion, distribute responsibility, and automate normativity. In doing so, they reconfigure not only what is done, but what can be imagined as possible, permissible, or desirable within the social field.

### v3

@@ Abstract
This article develops the concept of authoritarian recursion to theorize how artificial intelligence (AI) systems consolidate institutional control across education, military operations, and digital discourse. Rather than treating these domains in isolation, it identifies a shared recursive architecture in which algorithmic systems mediate judgment, obscure accountability, and reshape the conditions of moral and epistemic agency.
Grounded in critical discourse analysis and sociotechnical ethics, the paper synthesizes historical precedent, cultural narrative, and contemporary deployment to examine how intelligent systems normalize hierarchy under the guise of efficiency and neutrality. Case studies include automated proctoring in education, autonomous targeting in warfare, and algorithmic curation on social platforms. Cultural imaginaries such as Orwell’s Nineteen Eighty-Four, The Terminator’s Skynet, and Black Mirror are treated as heuristic devices that illuminate public anxieties and design assumptions embedded in technological systems.
The analysis integrates frameworks from the Fairness, Accountability, and Transparency (FAccT) paradigm, relational ethics, and data justice theory to explore the normative implications of predictive infrastructures. It argues that recursive control operates through moral outsourcing, behavioral normalization, and epistemic closure. By reframing AI not as a neutral tool but as a communicative and institutional infrastructure, the article highlights the need for ethical orientations that prioritize democratic refusal, epistemic plurality, and responsible design in the governance of intelligent systems.
Keywords: authoritarian recursion, algorithmic governance, recursive control, predictive infrastructures, sociotechnical ethics, platform power, educational surveillance, autonomous weapons, content moderation, critical discourse analysis, data justice, relational responsibility, infrastructural opacity, epistemic closure, democratic refusal
1 Introduction
Artificial intelligence (AI) is no longer merely a domain of technical optimization; it increasingly operates as an infrastructure of communication and governance. From remote proctoring platforms in education to autonomous weapons and content moderation systems, AI technologies are central to how institutions classify, interpret, and act on information. These systems do not simply automate decisions—they shape epistemic boundaries, delegate judgment, and redistribute communicative agency. In doing so, they embed themselves within broader sociotechnical arrangements of power and legitimacy.
This paper examines AI as a recursive mode of governance: a system that learns from, intervenes in, and ultimately reshapes the very behaviors it observes. These feedback architectures operate not only through prediction and abstraction, but through communicative invisibility. As platforms refine recommendations, military systems automate targeting, and educational software models behavioral suspicion, recursive AI infrastructures begin to enact what Couldry and Mejias term “data colonialism”—the capture and reconfiguration of human life through extraction and abstraction [1]. They reflect Gillespie’s notion of “infrastructural power,” where visibility and participation are mediated through algorithmic protocols [2], and contribute to what van Dijck et al. describe as “platform governance”—a mode of rule exercised through technical standards and informational architectures [3].
To theorize these dynamics, the article introduces the concept of authoritarian recursion. This term describes the self-reinforcing loops through which intelligent systems encode, legitimize, and propagate control logics—frequently under the rhetorical cover of personalization, neutrality, or operational efficiency. These recursive infrastructures obscure responsibility, foreclose contestation, and deepen asymmetries between users, institutions, and machinic systems.
The argument proceeds through three case domains—education, warfare, and digital discourse—where recursive AI systems materialize different forms of delegated authority and normative closure. Drawing on critical discourse analysis, the study treats AI not simply as a tool, but as a communicative actor that shapes who is visible, what is knowable, and which actions are thinkable. Cultural imaginaries such as Black Mirror and The Terminator’s Skynet are incorporated not as evidence, but as critical heuristics that reflect and amplify public anxieties around machinic autonomy and recursive control.
In dialogue with critical media studies, surveillance scholarship, and sociotechnical critique [4, 5, 6], the article advances the study of algorithmic governance by foregrounding recursion as both a technical logic and an ideological formation. Rather than proposing universal solutions or narrow design principles, the paper advocates for renewed scrutiny of how AI systems enact communicative authority—structuring infrastructures of attention, accountability, and legitimacy across institutional fields.
@@ 2 Literature Review
This section synthesizes scholarship on how AI technologies function as instruments of sociotechnical control in three domains: military automation, educational surveillance, and digital discourse. These sectors are often treated as distinct, yet their AI applications share design logics—opacity, delegated authority, and recursive feedback—that normalize institutional power under the guise of optimization. This literature provides the groundwork for theorizing authoritarian recursion as a mode of governance embedded in platforms and predictive systems. Drawing from ethical frameworks, historical precedents, and critical media theory, the review integrates both empirical and conceptual contributions to illustrate how automated systems entrench and legitimize normative authority across contexts.
@@ 2.1 Military AI and the Automation of Violence
These concerns align with the Fairness, Accountability, and Transparency (FAccT) framework, which holds that fairness in automated decision-making must extend beyond output metrics to include contextual sensitivity, procedural redress, and meaningful oversight [8, 9]. In military AI, accountability often becomes nominal—tethered to protocols rather than substantive ethical reflection.
Historical precedents further complicate the notion of technological neutrality. During World War II, IBM’s punch card infrastructure was deployed by Nazi Germany for logistics and census operations—systems later used to facilitate genocide [10]. These tools exemplified a bureaucratic rationality disturbingly resonant with today’s algorithmic architectures. As Asaro argues, the distancing of lethal decisions from moral responsibility through automation represents a dangerous ethical shift [11].
Speculative fiction reinforces this critique. Narratives like The Terminator, RoboCop, and Black Mirror episodes offer dystopian imaginaries of militarized AI. According to Cave et al., “the future imagined in fiction is often realized not because it is inevitable, but because it is ideologically compatible with dominant institutions” [12, p. 75]. These cultural texts anticipate how automation narratives rationalize political authority and normalize autonomous violence.
@@ 2.2 Educational AI: Automation of Surveillance and Discipline
Here, too, the FAccT triad is often invoked as a remedy. However, operational deployments rarely meet its normative thresholds. Mittelstadt et al. emphasize that fairness requires attention to context and historical inequalities, not just statistical parity [8]. Student-users typically lack access to the internal logic of these systems and have little recourse to challenge their outputs. Procedural fairness is often undermined when algorithmic opacity becomes the mechanism of control itself.
Noble argues that “algorithmic decision systems often act as new instruments of racial and economic profiling” [14, p. 34]. This echoes Selwyn’s concerns that digital surveillance in education creates a system “in which suspicion is automated and dissent is pathologized” [15]. McMillan Cottom [2020] further expands this critique by arguing that digital systems enact racial capitalism through “obfuscation as privatization and exclusion by inclusion,” framing technologies as infrastructures of sociopolitical ordering rather than neutral tools of progress. Together, these perspectives underscore how automated educational technologies reproduce the logic of panoptic discipline, as theorized by Foucault, where constant surveillance internalizes conformity [17].
Moreover, the political values behind these technologies often go unquestioned. Gilliard and Selwyn contend that “continued adoption of proctoring technologies in public education exposes a fundamental clash of politics,” where commercial priorities of security and efficiency override pedagogical values of equity and trust [18, p. 197]. These systems operationalize discipline not through direct coercion, but through the automation of suspicion and reduction of students to behavioral data.
@@ 2.3 AI in Discourse and Propaganda: Curation as Control
Gillespie emphasizes that platforms are not neutral hosts but “custodians of public discourse” who shape access to visibility through inscrutable recommendation logics [2, p. 197]. Zuboff describes this shift as “instrumentarian power,” wherein behavior is not repressed but tuned through predictive analytics and behavioral nudging [6, p. 377]. While Zuboff’s critique highlights platform logics of behavioral modification, the recursive dimension explored here adds a temporal structure to algorithmic authority.
@@
Together, these patterns constitute what we term authoritarian recursion—a self-reinforcing cycle in which AI technologies encode, naturalize, and propagate control logics across domains. This concept finds parallel in Bahrami’s Algemony framework, which similarly identifies AI’s capacity to reshape power through human-AI interactions, particularly via narrative modulation. Where authoritarian recursion emphasizes the structural inevitability of control through recursive feedback, Algemony reveals the Janus-faced nature of this process: AI systems exhibit both hegemonic reinforcement through delegated agency and disruptive potential through their inherent instability (e.g., via generative counter-narratives or unpredictable hyper-personalization). Both frameworks converge in their diagnosis of AI’s epistemic closure—what Zuboff terms the "instrumentarian" capture of human experience—while diverging in their emphasis on either the systemic (authoritarian recursion) or discursive (Algemony) dimensions of control. Building on Hanna and Kazim’s dignitarian ethics, we further recognize how these recursive systems violate fundamental principles of human dignity by: (1) instrumentalizing individuals through opaque algorithmic delegation, (2) distorting true human needs via self-reinforcing classifications, and (3) eroding privacy through pervasive surveillance infrastructures—violations exemplified in our case studies of educational proctoring and military targeting systems.
This synthesis lays the foundation for the case study analysis that follows, which illustrates how intelligent systems materialize these recursive dynamics in real-world governance structures while demonstrating the tensions between structural determinism and agential unpredictability in AI-mediated power. The dignitarian perspective provides crucial normative grounding for evaluating these systems, particularly in assessing when human oversight must remain irreducible to prevent dignity violations. Yet as Roy-Stang and Davies demonstrate through their analysis of cognitive biases, even well-intentioned governance interventions may be undermined by perceptual vulnerabilities that authoritarian recursion exploits—a challenge requiring both technical safeguards and epistemic humility in AI policy design.
@@ 3 Methodology
This inquiry adopts a qualitative, interpretive approach grounded in critical discourse analysis (CDA). Rather than treating AI as a purely technical system, the analysis considers how algorithmic infrastructures participate in the production of social meaning, institutional legitimacy, and normative authority. The aim is not to produce statistical generalization, but to interrogate how language, design, and abstraction coalesce in systems that structure governance across multiple domains.
CDA provides a framework for understanding AI as both discursive and material. Following Fairclough, discourse is treated as “both constitutive and constituted”—a site where social relations are encoded, contested, and reproduced [25]. This duality makes CDA especially suited to analyzing recursive architectures that simultaneously reflect institutional norms and help reshape them through feedback and optimization.
The empirical and conceptual materials informing this analysis come from three interlinked domains. First, peer-reviewed academic literature in media studies, science and technology studies (STS), and AI ethics provides critical foundations for identifying the normative and infrastructural dimensions of algorithmic systems [8, 5, 26, 2]. Second, fictional narratives—such as Orwell’s Nineteen Eighty-Four, Black Mirror episodes, and The Terminator—are incorporated not as data points but as cultural diagnostics. These texts are mobilized heuristically to surface the imaginaries, anxieties, and anticipatory logics that often shape both public understanding and institutional development of AI systems [27, 28, 29]. Third, contemporary policy documents and technical implementations—including military doctrine on lethal autonomous weapons, algorithmic proctoring guidelines, and platform transparency reports—offer grounded illustrations of how recursive infrastructures operate across sectors [9, 15].
The analytical process unfolded in three stages. First, key texts and artifacts were situated within their institutional and historical contexts, with attention to their role in shaping or reflecting dominant discourses on automation, governance, and visibility. Second, themes were identified through iterative coding, blending normative frameworks such as Fairness, Accountability, and Transparency (FAccT) with emerging concerns around relational ethics, surveillance asymmetries, and predictive control. Finally, patterns were synthesized across the three case domains—education, warfare, and discourse—to trace how recursive structures propagate common design logics and ethical blind spots.
The interpretive stance taken here is explicitly critical. Rather than assuming AI systems are neutral or inevitable, the analysis foregrounds the political and ideological work performed by these infrastructures. While the scope is illustrative rather than exhaustive, the goal is to render visible how recursion operates not only as a computational feature but as a sociotechnical logic of governance—shaping what can be known, predicted, and contested in the age of algorithmic mediation.
@@ 4 Case Studies
AI systems are increasingly embedded in everyday infrastructures, where they mediate visibility, classification, and decision-making. This section examines how recursive architectures of control materialize across three institutional domains—education, warfare, and digital discourse. While these cases are sectorally distinct, they converge in their design assumptions: opacity, delegation of judgment, and feedback-driven prediction. The aim is not to compare implementations at the technical level but to trace how these systems encode normative authority while displacing accountability.
4.1 Education: Algorithmic Proctoring and the Automation of Suspicion
Remote proctoring platforms such as Proctorio and ExamSoft exemplify how AI systems reshape pedagogical environments under the guise of academic integrity. These systems rely on webcam feeds, gaze detection, keystroke tracking, and ambient audio to flag “suspicious” behavior. Yet these classifications are rarely transparent, and often correlate with racialized and gendered biases in facial recognition technology. Williams et al. show that students with darker skin tones—particularly Black students and women of color—are disproportionately flagged, not due to misconduct, but due to technical misrecognition [13].
The act of learning becomes a site of surveillance, and intellectual agency is restructured into a performance of algorithmic legibility. As Gilliard and Selwyn argue, such systems reflect a “clash of politics,” where commercial logics of risk management and datafication override pedagogical values such as trust, dialogue, and equity [18]. These dynamics reproduce what Foucault identified as disciplinary power, where visibility serves as a mode of behavioral regulation [17].
While appeals to FAccT principles may offer procedural remedies, the underlying infrastructure of suspicion remains intact. Students often lack access to the reasoning behind automated flags and have limited recourse for contestation. In effect, these systems reframe education not as relational learning, but as a risk environment managed through probabilistic surveillance.
4.2 Warfare: Autonomous Systems and Delegated Lethality
The military deployment of AI systems—from Israel’s Harpy drones to the U.S. MQ-9 Reaper—represents a deepening entanglement between automation and coercive force. These systems promise speed and precision, yet shift ethical deliberation into technical procedure. As Asaro warns, the automation of lethal decision-making distances human agents from moral responsibility and erodes existing norms under international humanitarian law [11].
Fictional narratives such as *Skynet* in *The Terminator* resonate here not as prophecy, but as cautionary heuristics. Skynet’s recursive logic—learning from threat environments, optimizing for efficiency, and ultimately bypassing human oversight—reflects a broader pattern in military automation: operational assumptions become epistemic truths when encoded into training data and model outputs. These systems increasingly define what counts as a threat based on prior engagements, creating closed loops of risk perception and response.
Cave and Dihal argue that such narratives shape public discourse and institutional choices by making visible the ideological compatibility between automation and governance [27]. Far from exaggerating reality, speculative fiction reveals the recursive rationality that guides contemporary weapon systems—logics that prioritize abstraction and autonomy while suppressing ambiguity and contestation.
4.3 Discourse: Algorithmic Curation and the Infrastructural Shaping of Visibility
In digital media environments, algorithmic curation systems govern what users see, believe, and share. Platforms like Facebook, YouTube, and TikTok rank and recommend content based on engagement metrics, which recursively train the algorithms to maximize attention capture. This logic not only polarizes discourse but normalizes visibility as a function of behavioral prediction.
Gillespie describes these systems as “custodians of public discourse,” whose infrastructural design choices shape the epistemic conditions of civic life [2]. Zuboff’s critique of “instrumentarian power” emphasizes that such platforms do not merely moderate content—they model and nudge behavior at scale through continuous feedback loops [6]. These loops, once set in motion, generate their own reality: prior interactions inform future recommendations, entrenching echo chambers and limiting epistemic diversity.
Interventions premised on FAccT often fall short, as transparency mechanisms rarely expose the full contours of algorithmic influence. As Mittelstadt et al. argue, ethical frameworks lose traction when deployed as procedural checklists without structural transformation [8]. Instead of accountability, platforms offer explainability, often in the form of generic disclosures that do little to empower users or regulators.
The recursive character of algorithmic discourse governance mirrors historical propaganda infrastructures in its effects, if not its form. Where authoritarian regimes once synchronized opinion through mass spectacle, platforms now achieve similar outcomes through personalization and behavioral profiling. The result is not ideological uniformity, but a segmented public sphere optimized for affective intensification and informational control [20, 19].
These cases reveal how AI systems do not merely optimize existing functions—they reconstitute the meaning of legitimacy, trust, and agency. The recursive architectures examined here are not errors of implementation, but expressions of a broader political rationality: one in which prediction supplants deliberation, abstraction substitutes for context, and visibility becomes a privilege granted by infrastructural design. As the following analysis will show, these recursive forms of control demand not only technical reform, but structural and discursive interrogation.